Numerical Analysis I, Fall 2017 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 11

1. Section 4.4: Problems 7, 14, 23, 24, 26. Problem 7 is a programming assignment. Make good use of the ’for’ loop in matlab/C, or ’sum’ in matlab.

2. Section 4.7: Problems 2(b), 4(b), 11, 13, 14.